Grammars and Lexicons: 11-721 Fall Term, 2003 Extra Credit Homework

1 Due Date

You can hand in extra credit homework any time before November 1.

2 Grading

We have not yet decided how the extra credit homework will affect your grade.

3 The Assignment: Linguistlist Discussion Board

The Linguistlist (www.linguistlist.org) has a discussion board for interesting and controversial topics. The discussion board is at: http://linguistlist.org/issues/indices/Disc2003r.html. Do the following for extra credit:

• Choose a discussion topic. The topics I recommend are:

– NYT Essay on Endangered Languages

– BBC Story: Mandarin and Brain Hemispheres

– Celtic found to have ancient roots

– Anaphora makes the headlines

Don’t choose “Academic Boycotts”. If you want to choose another topic (other than the ones listed above), please check with me first.

• If you choose the NYT or BBC topics, find the original article and read it. (You don’t have to find the original Celtic article; one of the reviewers says that it is too hard to read.) I didn’t see a pointer to the original Anaphora article, but see if you can find it by searching on the web.

• Read all the messages in the thread of the topic you chose.

• Summarize each point that is made in the thread or the original article. If two people make the same point, you can combine them.

• For each point that you summarize, write in bold face “CONVINCED” or “NOT CONVINCED”.

• Write your own summary (less than one page) of your opinion of the issue.

• Give me a draft to review and comment on before handing in the final version.